# Constraining Power of Wavelet vs. Power Spectrum Statistics for CMB Lensing and Weak Lensing with Learned Binning

### v1

@@ Abstract
We present forecasts for constraints on the matter density ($\Omega_{m}$) and the amplitude of matter density fluctuations at 8h-1Mpc ($\sigma_{8}$) from CMB lensing convergence ($\kappa_{\text{CMB}}$) maps and galaxy weak lensing convergence ($\kappa_{\text{WL}}$) maps. For $\kappa_{\text{CMB}}$ auto statistics, we compare the angular power spectra ($C_{\ell}$’s) to the wavelet scattering transform (WST) coefficients. For $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ statistics, we compare the cross angular power spectra to wavelet phase harmonics (WPH). This work also serves as the first application of WST and WPH to these probes. For $\kappa_{\text{CMB}}$, we find that WST and $C_{\ell}$’s yield similar constraints in forecasts for the Simons Observatory and the South Pole Telescope. However, WST gives a tighter constraint on $\sigma_{8}$ by a factor of $1.7$ for Planck data. When $\kappa_{\text{CMB}}$ is crossed with $\kappa_{\text{WL}}$ projected from Euclid Data Release 2 (DR2), we find that WPH outperforms cross-$C_{\ell}$’s by factors between $2.4$ and $3.8$ for individual parameter constraints. To compare these different summary statistics we develop a novel learned binning approach. This method compresses summary statistics while maintaining interpretability. We find this leads to improved constraints compared to more naive binning schemes for $C_{\ell}$’s, WST, and most significantly WPH. By learning the binning and measuring constraints on distinct data sets, our method is robust to overfitting by construction.
[…]
0.75in
1 Introduction
@@
Gravitational lensing of photons occurs when light is deflected by a mass distribution in between the source and the observer (Einstein, 1936). By observing distortions in the shapes of numerous galaxies from weak lensing (hereafter WL), statistical methods can be used to obtain integrated line-of-sight mass maps (e.g. Chang et al., 2015). Current and upcoming surveys, such as the Euclid mission (Bretonnière et al., 2023), the Dark Energy Survey (DES, Collaboration, 2016), the Kilo-Degree Survey (KIDS, Kuijken et al., 2015), the Hyper Suprime-Cam (HSC, Aihara et al., 2017), and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST, Collaboration et al., 2009), will measure galaxy shapes, enabling measurements of these integrated mass maps. Additionally, by observing the lensing of the cosmic microwave background (hereafter CMBL), one can obtain mass maps integrated all the way back to the surface of last scattering (e.g. Madhavacheril et al., 2023). Examples of surveys that perform this measurement include Planck (Aghanim et al., 2020), the Atacama Cosmology Telescope (ACT) (Aiola et al., 2020), the South Pole Telescope (SPT) (Simard et al., 2018), and the Simons Observatory (SO) (Ade et al., 2019). Lensing-convergence maps from these surveys are used to constrain cosmology by comparing measured summary statistics with predictions from analytic theories or forward-modeled simulations, within either a likelihood-based or simulation-based inference framework.
@@
If the convergence maps were Gaussian random fields, all of their information would be captured by two-point statistics. However, due to non-linear gravitational evolution, the large-scale structure of the universe has non-Gaussianities that become more pronounced at low redshift. Since two-point statistics fail to capture non-Gaussian information, other summary statistics are desirable. Numerous summary statistics have been used to try to extract non-Gaussian information, including the wavelet scattering transform (e.g. Cheng et al., 2020; Valogiannis & Dvorkin, 2022a, b; Valogiannis et al., 2024b; Gatti et al., 2023; Cheng et al., 2025) and wavelet phase harmonics (e.g. Gatti et al., 2023). Other approaches include higher-order N-point statistics (e.g. Gil-Marín et al., 2015; Philcox et al., 2021; Ivanov et al., 2022; Chudaykin et al., 2025; Sugiyama et al., 2025), density-split clustering (Paillas et al., 2024), topological descriptors such as Minkowski functionals (Liu et al., 2025; Prat et al., 2025), and machine-learning-based summaries like convolutional neural networks (Jeffrey et al., 2025).
@@
In this work, we compare the information content of selected summary statistics applied to the lensing of the cosmic microwave background and weak lensing fields using Fisher forecasts. Each summary statistic compresses the information in a field and typically depends on a handful of hyper-parameters (such as the range of scales in a power spectrum or the number of different opening angles in three-point statistics). Hyperparameter choices change the effective scales and the dimensionality of the summary statistics; this can impact the stability of the covariance inversion in Gaussian-likelihood pipelines and, for simulation-based inference pipelines, inflate model complexity and training variance. Constraints become then highly dependent on hyper-parameters. We address this by proposing and developing a novel learned binning approach, which produces compact and informative summaries.
@@
The rest of the paper is organized as follows: in Section 2 we introduce the simulations used in our analysis and explain how to use them to construct convergence maps. In Section 3, we introduce the summary statistics used. In Section 4, we present an overview of our learned binning approach, which is then applied to simulated data in Section 5. Finally, we conclude in Section 6. All parameter constraints are presented in Appendix A, and their convergence is tested in Appendix B. Additional details on the learned binning approach are given in Appendix C. We verify the Gaussianity of our summary statistics required for a reliable Fisher forecast in Appendix D.
2 Simulated Observables
@@
where $D_{L}$, $D_{S}$, and $D_{LS}$ are the angular distances from the observer to the lens, the observer to the source, and the lens to the source respectively, $c$ is the speed of light, and $\vec{\theta}$ is the angular position on the sky. From the lensing potential, we can then construct the convergence map $\kappa$ as:
@@
2.1 Simulations: ULAGAM Suite
The analysis in this work uses the ULAGAM suite of N-body simulations (Anbajagane et al., 2024). The suite consists of 2000 simulations at a fiducial cosmology, plus sets of 100 simulations where one parameter is varied at a time in a stepwise fashion for Fisher forecast purposes.
@@ 2.2 CMB Lensing Convergence
Figure 1: We show the average noiseless auto power spectrum of $\kappa_{\text{CMB}}$ (black) in comparison to noise power spectra from all CMB surveys used (dashed lines). For ACT, we only plot until $\ell=2090$ as this is the limit of our noise power spectrum data. For SPT, we plot the noise power spectra from the main and summer fields separately. The cusps (for instance at $\ell\approx 2200$ for SPT) occur because our noise power spectra are linear interpolations between specified points.
@@
We follow the procedure described in Section $3.1$ of Omori (2022) to create noiseless, full-sky $\kappa_{\text{CMB}}$ maps for each of our simulations. In particular, we use the Born approximation to integrate the overdensity maps up to redshift $z=3.5$. The CMB lensing power spectrum is then computed for the range $3.5<z<1089$, and a Gaussian realization based on this spectrum is generated and added to obtain the full $\kappa_{\text{CMB}}$ map.
@@
We generate mock survey $\kappa_{\text{CMB}}$ maps by adding Gaussian noise and footprint masks to the full-sky, noiseless convergence maps, simulating various experiments. Specifically, we produce maps for Planck, ACT Data Release 6 (DR6), SPT main and summer fields, and SO. The noise is modeled as purely Gaussian and generated from the corresponding noise power spectra, taken from Ade et al. (2019) and Prabhu et al. (2024). For ACT, our noise power spectra only go to $\ell_{\text{max}}=2090$, so we set all signal beyond this to zero. Signal and noise power spectra for $\kappa_{\text{CMB}}$ are shown for all surveys used in Figure 1.
@@
For Planck, we further apply a Gaussian smoothing beam with a full width half maximum of 10 arcminutes after adding noise. This mimics the instrument’s beam. We neglect the beam of the other experiments since they are smaller than the pixel size of our maps. Survey masks are applied to each simulated map to match the respective sky coverage. For SPT, we consider the union of the main and summer fields, each region with its corresponding noise power spectrum applied. The coverage area of all surveys can be found in Table 1. An example $\kappa_{\text{CMB}}$ for Planck levels of noise is shown in Figure 2.
2.3 Weak Galaxy Lensing Convergence
@@
To create mock $\kappa_{\text{WL}}$ maps, we use a simplified version of the pipeline described in Gatti et al. (2023), tailored to resemble Euclid DR2 data. First, we generate noiseless convergence maps as a function of redshift from the density field using the Born approximation.
@@
We assume that the Euclid DR2 weak-lensing source sample is divided into six tomographic bins of roughly equal number density. For simplicity, we only use two tomographic bins: the bin with the lowest redshift and the bin with the highest redshift. The first (last) tomographic bin has a mean redshift of 0.33 (1.37) and a standard deviation of 0.09 (0.16). For each bin, we produce a convergence map by summing the redshift shells weighted by the corresponding redshift distribution. We add Gaussian shape noise, assuming an ellipticity dispersion of $\sigma_{e}=0.26$ and a galaxy number density of 30 galaxies per arcmin2, uniformly distributed across the tomographic bins (Bretonnière et al., 2023). Finally, we mask the mock $\kappa_{\text{WL}}$ maps by applying the Euclid DR2 footprint mask. For all $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ statistics, we apply both the Euclid and CMB-survey masks. The overlap coverage area for all surveys can be found in Table 1.
3 Summary Statistics
@@ 3.1 Power Spectrum
The auto angular power spectra are defined with $A=B$. From here on, we will omit superscripts as the context will make it clear which power spectrum we are referring to.
@@
3.2 Wavelet Scattering Transform
The WST was initially proposed in Mallat (2012) and then used in cosmology for extracting information from weak lensing maps (Cheng et al., 2020) and galaxy clustering (Valogiannis & Dvorkin, 2022b, a; Valogiannis et al., 2024b). WST consists of convolving the initial map with a directional, multi-scaled wavelet and taking the modulus of the result. Multiple successive convolutions can be performed to probe non-Gaussianities. An average over all pixels is used to convert any map into a scattering coefficient, the building block of this summary statistic.
@@ 3.2 Wavelet Scattering Transform
In the above, $\psi$ is the localized wavelet. In general, $\psi$ will have a scale and a directional dependence, represented here by the integer-valued parameters $j$ and $l$ respectively. In this case, one can obtain the first three orders of scattering coefficients as:
@@
where $\langle$$\rangle$ is a spatial average. As in other works (e.g. Gatti et al., 2023; Cheng et al., 2025), we only use $S_{1}$ and $S_{2}$ coefficients. If the overlap of wavelets $\psi_{j,l}$ and $\psi_{j,l+1}$ is small and we have statistical isotropy, we can average over orientations without losing information: $S_{1}(j_{1},l_{1})\rightarrow S_{1}(j_{1})$ and $S_{2}(j_{1},l_{1},j_{2},l_{2})\rightarrow S_{2}(j_{1},j_{2},|l_{1}-l_{2}|)$. In this work we assume isotropy for all relevant summary statistics and average over opening angles. Additionally, not all scale combinations contribute to $S_{2}$. As shown in Cheng & Ménard (2021), only $S_{2}$ coefficients with $j_{2}>j_{1}$ are significant.
@@
where $\sigma=0.8\times 2^{j}$, $k_{0}=0.75\pi\times 2^{-j}$, $\text{arg}(\vec{k}_{0})=\pi\theta/L$, with $\theta=(\frac{L}{2}-1-l)\pi/L$. The counterterm $\beta=\exp\left(-\frac{k_{0}^{2}\sigma^{2}}{2}\right)$ is introduced so that $\tilde{\psi}(0)=0$. The $j$ index will take integer values between 0 and J1J-1, where $J$ controls the number of scales and is at most $\log_{2}(\text{Image Size})-1$. This keeps the wavelet peaks in Fourier space roughly between the fundamental frequency and the Nyquist frequency of the image. The $l$ values range between 0 and L1L-1, where $L$ specifies the number of different angles to probe. While Morlet wavelets traditionally have the above dyadic spacing, we also experiment with finer spacing. Details on this will be given explicitly before any results are shown.
To apply the wavelet scattering transform, the input image must be two dimensional and flat. To achieve this with spherical data we split the data into 768 disjoint patches, which are the HEALPix pixels at $\texttt{NSIDE}=8$. The center of each of these patches is rotated to the north pole of the sphere. From here, gnomonic projection is used to project the high resolution patch onto a $256\times 256$ grid (as in Gatti et al., 2023), which sets our maximum $J$ value at $J=7$. Due to spherical distortions, not all HEALPix pixels are mapped one-to-one onto a pixel in the final grid. We do not attempt to correct for this effect in this work. After getting scattering coefficients for each patch, we average across patches to get scattering coefficients for the entire map.
@@
3.3 Wavelet Phase Harmonics

### v2

@@ Abstract
We present forecasts for constraints on the matter density ($\Omega_{m}$) and the amplitude of matter density fluctuations at 8h-1Mpc ($\sigma_{8}$) from CMB lensing convergence ($\kappa_{\text{CMB}}$) maps and galaxy weak lensing convergence ($\kappa_{\text{WL}}$) maps. For $\kappa_{\text{CMB}}$ auto statistics, we compare the angular power spectra ($C_{\ell}$’s) to the wavelet scattering transform (WST) coefficients. For $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ statistics, we compare the cross angular power spectra to wavelet phase harmonics (WPH). This work also serves as the first application of WST and WPH to these probes. For $\kappa_{\text{CMB}}$, we find that WST and $C_{\ell}$’s yield similar constraints in forecasts for all surveys considered in this work. When $\kappa_{\text{CMB}}$ is crossed with $\kappa_{\text{WL}}$ projected from Euclid Data Release 2 (DR2), we find that WPH outperforms cross-$C_{\ell}$’s by factors between $2.2$ and $3.4$ for individual parameter constraints. To compare these different summary statistics, we develop a novel learned binning approach. This method compresses summary statistics while maintaining interpretability. We find this leads to improved constraints compared to more naive binning schemes for our wavelet-based statistics, but not for $C_{\ell}$’s. By learning the binning and measuring constraints on distinct data sets, our method is robust to overfitting by construction.
I Introduction
@@
Gravitational lensing of photons occurs when light is deflected by a mass distribution between the source and the observer [16]. By observing distortions in the shapes of numerous galaxies from weak lensing (hereafter WL), statistical methods can be used to obtain integrated line-of-sight mass maps [e.g. 8]. Current and upcoming surveys, such as the Euclid mission [7], the Dark Energy Survey [DES, 14], the Kilo-Degree Survey [KIDS, 26], the Hyper Suprime-Cam [HSC, 3], and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time [LSST, 13], will measure galaxy shapes, enabling measurements of these integrated mass maps. Additionally, by observing the lensing of the cosmic microwave background (hereafter CMBL), one can obtain mass maps integrated all the way back to the surface of last scattering [e.g. 29]. Examples of surveys that perform this measurement include Planck [2], the Atacama Cosmology Telescope (ACT) [4], the South Pole Telescope (SPT) [42], and the Simons Observatory (SO) [1]. Lensing-convergence maps from these surveys are used to constrain cosmology by comparing measured summary statistics with predictions from analytic theories or forward-modeled simulations, within either a likelihood-based or simulation-based inference framework.
@@
If the convergence maps were Gaussian random fields, all of their information would be captured by two-point statistics. However, due to non-linear gravitational evolution, the large-scale structure of the universe has non-Gaussianities that become more pronounced at low redshift. Since two-point statistics fail to capture non-Gaussian information, other summary statistics are desirable. Numerous summary statistics have been used to try to extract non-Gaussian information, including the wavelet scattering transform [e.g. 11, 46, 47, 49, 18, 9] and wavelet phase harmonics [e.g. 18]. Other approaches include higher-order N-point statistics [e.g. 19, 36, 24, 12, 43], density-split clustering [34], topological descriptors such as Minkowski functionals [28, 39, 6], and machine-learning-based summaries like convolutional neural networks [17, 25].
@@
In this work, we compare the information content of selected summary statistics applied to the lensing of the cosmic microwave background and weak lensing fields using Fisher forecasts. Each summary statistic compresses the information in a field and typically depends on a handful of hyperparameters (such as the range of scales in a power spectrum or the number of different opening angles in three-point statistics). Hyperparameter choices change the effective scales and the dimensionality of the summary statistics; this can impact the stability of the covariance inversion in Gaussian-likelihood pipelines and, for simulation-based inference pipelines, inflate model complexity and training variance. Constraints become then highly dependent on hyperparameters. We address this by proposing and developing a novel learned binning approach, which produces compact and informative summaries.
@@
The rest of the paper is organized as follows: in Section II we introduce the simulations used in our analysis and explain how to use them to construct convergence maps. In Section III, we introduce the summary statistics used. In Section IV, we present an overview of our learned binning approach, which is then applied to simulated data in Section V. Finally, we conclude in Section VI. All parameter constraints are presented in Appendix A, and their convergence is tested in Appendix B. We test the convergence of our derivatives in Appendix C. Additional details on the learned binning approach are given in Appendix D. We verify the Gaussianity of our summary statistics required for a reliable Fisher forecast in Appendix E.
II Simulated Observables
@@
where $D_{L}$, $D_{S}$, and $D_{LS}$ are the angular distances from the observer to the lens, the observer to the source, and the lens to the source, respectively, $c$ is the speed of light, and $\vec{\theta}$ is the angular position on the sky. From the lensing potential, we can then construct the convergence map $\kappa$ as:
@@
II.1 Simulations: ULAGAM Suite
The analysis in this work uses the ULAGAM suite of N-body simulations [5]. The suite consists of 2000 simulations at a fiducial cosmology, plus sets of 100 simulations where one parameter is varied at a time in a stepwise fashion for Fisher forecast purposes.
@@ II.2 CMB Lensing Convergence
Figure 1: We show the average noiseless auto power spectrum of $\kappa_{\text{CMB}}$ (black) in comparison to noise power spectra from all CMB surveys used (dashed lines). For ACT, we only plot until $\ell=2090$ as this is the limit of our noise power spectrum data. For SPT, we plot the noise power spectra from the main and summer fields separately. The cusps (for instance, at $\ell\approx 2200$ for SPT) occur because our noise power spectra are linear interpolations between specified points.
@@
We follow the procedure described in Section $3.1$ of Omori [33] to create noiseless, full-sky $\kappa_{\text{CMB}}$ maps for each of our simulations. In particular, we use the Born approximation to integrate the overdensity maps up to redshift $z=3.5$. The CMB lensing power spectrum is then computed for the range $3.5<z<1089$, and a Gaussian realization based on this spectrum is generated and added to obtain the full $\kappa_{\text{CMB}}$ map.
@@
We generate mock survey $\kappa_{\text{CMB}}$ maps by adding Gaussian noise and footprint masks to the full-sky, noiseless convergence maps, simulating various experiments. Specifically, we produce maps for Planck, ACT Data Release 6 (DR6), SPT main and summer fields, and SO. The noise is modeled as purely Gaussian and generated from the corresponding noise power spectra, taken from Ade et al. [1] and Prabhu et al. [38]. For ACT, our noise power spectra only go to $\ell_{\text{max}}=2090$, so we set all signal beyond this to zero. Signal and noise power spectra for $\kappa_{\text{CMB}}$ are shown for all surveys used in Figure 1.
@@
For Planck, we further apply a Gaussian smoothing beam with a full width half maximum of 10 arcminutes after adding noise. This mimics the instrument’s beam. We ignore the beams of the other experiments because they are smaller than the pixel size of our maps. Survey masks are applied to each simulated map to match the respective sky coverage. For SPT, we consider the union of the main and summer fields, each region with its corresponding noise power spectrum applied. The coverage area of all surveys can be found in Table 1. An example $\kappa_{\text{CMB}}$ for Planck levels of noise is shown in Figure 2.
We note that this procedure is idealized and produces simplified CMB lensing maps with Gaussian noise and no additional contaminants. These maps are sufficient for the purpose of this work, namely, a forecast-level assessment of the intrinsic non-Gaussian information in the lensing field, but they are not suitable for direct comparison with real data. In practice, CMB lensing is not measured directly but reconstructed from multi-frequency temperature and polarization maps using, for example, quadratic estimators [23], after applying standard data-cleaning steps such as frequency filtering and survey-mask corrections. Since lensing is a small signal superimposed on the primary CMB anisotropies, the reconstructed maps are contaminated by the CMB itself and by reconstruction effects that must be calibrated using simulations. Realistic CMB lensing mocks, therefore, require simulating unlensed CMB maps, lensing them with a full-sky convergence field, adding realistic anisotropic and non-Gaussian noise, and reconstructing the lensing field using the same estimators and corrections applied to the data [27]. The resulting maps naturally include non-Gaussian reconstruction noise and higher-order bias terms. Our aim here is more limited: to quantify, at the forecast level, the amount of intrinsic non-Gaussian information in the lensing field itself. For this purpose, modelling the noise as Gaussian is sufficient to capture the correct order-of-magnitude behavior, as we do not attempt to reproduce the detailed statistics of reconstructed CMB lensing maps.
II.3 Weak Galaxy Lensing Convergence
@@
To create mock $\kappa_{\text{WL}}$ maps, we use a simplified version of the pipeline described in Gatti et al. [18], tailored to resemble Euclid DR2 data. First, we generate noiseless convergence maps as a function of redshift from the density field using the Born approximation.
@@
We assume that the Euclid DR2 weak-lensing source sample is divided into six tomographic bins of roughly equal number density. For simplicity, we use only two tomographic bins: the lowest-redshift bin and the highest-redshift bin. The first (last) tomographic bin has a mean redshift of 0.33 (1.37) and a standard deviation of 0.09 (0.16). For each bin, we produce a convergence map by summing the redshift shells weighted by the corresponding redshift distribution. We add Gaussian shape noise, assuming an ellipticity dispersion of $\sigma_{e}=0.26$ and a galaxy number density of 30 galaxies per arcmin2, uniformly distributed across the tomographic bins [7]. Finally, we mask the mock $\kappa_{\text{WL}}$ maps by applying the Euclid DR2 footprint mask. For all $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ statistics, we apply both the Euclid and CMB-survey masks. The overlap coverage area for all surveys can be found in Table 1.
III Summary Statistics
@@ III.1 Power Spectrum
The auto angular power spectra are defined with $A=B$. From here on, we will omit superscripts as the context will make it clear which power spectrum we are referring to. The $C_{\ell}$’s are all computed on the masked fields described in Section II.
@@
III.2 Wavelet Scattering Transform
The WST was initially proposed in Mallat [30] and then used in cosmology to reduce noise in simulations [41] and extract information from weak lensing maps [11] and galaxy clustering [47, 46, 49]. WST consists of convolving the initial map with a directional, multi-scaled wavelet and taking the modulus of the result. Multiple successive convolutions can be performed to probe non-Gaussianities. An average over all pixels is used to convert any map into a scattering coefficient, the building block of this summary statistic.
@@ III.2 Wavelet Scattering Transform
In the above, $\psi$ is the localized wavelet. In general, $\psi$ will have a scale and a directional dependence, represented here by the integer-valued parameters $j$ and $l$, respectively. In this case, one can obtain the first three orders of scattering coefficients as:
@@
where $\langle$$\rangle$ is a spatial average. As in other works [e.g. 18, 9], we only use $S_{1}$ and $S_{2}$ coefficients. If the overlap of wavelets $\psi_{j,l}$ and $\psi_{j,l+1}$ is small and we have statistical isotropy, we can average over orientations without losing information: $S_{1}(j_{1},l_{1})\rightarrow S_{1}(j_{1})$ and $S_{2}(j_{1},l_{1},j_{2},l_{2})\rightarrow S_{2}(j_{1},j_{2},|l_{1}-l_{2}|)$. In this work, we assume isotropy for all relevant summary statistics and average over opening angles. Additionally, not all scale combinations contribute to $S_{2}$. As shown in Cheng and Ménard [10], only $S_{2}$ coefficients with $j_{2}>j_{1}$ are significant.
@@
where $\sigma=0.8\times 2^{j}$, $k_{0}=0.75\pi\times 2^{-j}$, $\text{arg}(\vec{k}_{0})=\theta$, with $\theta=(\frac{L}{2}-1-l)\pi/L$. The counterterm $\beta=\exp\left(-\frac{k_{0}^{2}\sigma^{2}}{2}\right)$ is introduced so that $\tilde{\psi}(0)=0$. The $j$ index will take integer values between 0 and J1J-1, where $J$ controls the number of scales and is at most $\log_{2}(\text{Image Size})-1$. This keeps the wavelet peaks in Fourier space roughly between the fundamental frequency and the Nyquist frequency of the image. The $l$ values range between 0 and L1L-1, where $L$ specifies the number of different angles to probe. While Morlet wavelets traditionally have the above dyadic spacing, we also experiment with finer spacing. Details on this will be given explicitly before any results are shown.
To apply the wavelet scattering transform, the input image must be two-dimensional and flat. To achieve this with spherical data, we split the data into 768 disjoint patches, which are the HEALPix pixels at $\texttt{NSIDE}=8$. The center of each of these patches is rotated to the north pole of the sphere. From here, gnomonic projection is used to project the high-resolution patch onto a $256\times 256$ grid [as in 18], which sets our maximum $J$ value at $J=7$. Due to spherical distortions, not all HEALPix pixels are mapped one-to-one onto a pixel in the final grid. We do not attempt to correct for this effect in this work. After obtaining scattering coefficients for each patch, we average them to obtain scattering coefficients for the entire map.
@@
III.3 Wavelet Phase Harmonics